OpenStudy (anonymous):

The diameter of a circular flower bed is 10 feet. Estimate the circumference of the flower bed.

OpenStudy (anonymous):

radius=diameter/2 r=5ft circumference=2\[\pi \]r =2*5*\[\pi \] =10\[\pi \]

OpenStudy (anonymous):

radius= diameter / 2 radius=5ft circumference = 2*pi*r = 2*5*pi =10pi ft